A bond with a par value of $100 is currently trading at a price of $104....

  1. A bond with a par value of $100 is currently trading at a price of $104. The bond has a coupon rate of 4.16%, and it has exactly 10 years remaining until maturity. What is the bond’s current yield?

  2. An investor buys a bond with a $100 par value and a 5% coupon rate for $97. The bond pays interest semiannually. Exactly one year later, just after receiving the second coupon payment, the investor sells the bond for $96. What was the investor’s rate of return over the year from owning the bond?
Answer #1

1.

NPER 20 [10*2]
FV 100
PMT 4.16 [1000*4.16%]
PV 104
Current yield 3.87% [Rate ( nper, pmt,-pv,fv)]

2. Coupon payment = 100* 5% = $5

Loss on payment = $97 - $ 96 = $1

Return = $5 - $1 = $4

Rate of return = ($4 / $97) *100 = 4.123%